# Energy Extraction via Magnetic Reconnection in Konoplya-Rezzolla-Zhidenko Parameterized Black Holes

### v1

@@ Abstract
Recently, Comisso and Asenjo proposed a novel mechanism for harnessing energy from black holes through magnetic reconnection. Our study focuses on exploring the utilization of this mechanism on Konoplya-Rezzolla-Zhidenko (KRZ) parameterized black holes to assess the impact of deformation parameters on energy extraction. Among the various parameters, $\{\delta_{1},\delta_{2}\}$ are identified as the most important factors affecting the physics under consideration. The influence of these two parameters on the ergoregion, circular geodesics in the equatorial plane, and energy extraction from the KRZ black holes through magnetic reconnection is carefully analyzed. Results indicate that deviations from the Kerr metric have a notable influence on the energy extraction process. Particularly, energy extraction is enhanced with more negative ${\delta_{1}}$ or more positive ${\delta_{2}}$ within the range constrained by current astronomical observations, resulting in significantly higher maximum power and efficiency compared to the Kerr model. Moreover, when ${\delta_{2}}$ is sufficiently negative, extracting energy through this mechanism becomes increasingly challenging, necessitating an exceptionally high black hole spin.
@@ I Introduction
The first plausible mechanism for energy extraction from black holes was proposed by Penrose in 1969, now known as the Penrose process [8]. The Penrose process involves the splitting of a particle in the ergoregion into two particles. Measured from infinity, one of these particles may carry negative energy, while the other carries excess positive energy. The extraction of energy from a black hole occurs when the negative energy particle is absorbed by the black hole and the positive energy particle escapes. However, the Penrose process is considered impractical because it requires the two newborn particles to have relative velocities greater than 1/2 the speed of light, an event thought to be rare in real astrophysical processes [9, 10]. Inspired by Penrose’s work, physicists have proposed several alternative mechanisms for harnessing energy from black holes, including superradiant scattering [11], the collisional Penrose process [12], the Blandford-Znajek (BZ) process [13], and the magnetohydrodynamic (MHD) Penrose process [14]. Among these mechanisms, the BZ process is currently regarded as the most promising for elucidating gamma-ray bursts (GRBs) [15, 16, 17] and relativistic jets in active galactic nuclei (AGNs) [18, 19, 20, 21].
@@
Recently, based on an exploratory study [22], a novel mechanism for black hole energy extraction has been proposed by Comisso and Asenjo [23]. In this mechanism, energy extraction is realized by a process of magnetic reconnection in the ergoregion of a rotating black hole. The process begins with the magnetic field around the black hole forming anti-parallel configurations near the equatorial plane due to the black hole’s rotation [24, 25, 26, 27]. The jumps in the magnetic field direction induce the formation of current sheets between them, which are destroyed by plasmoid instability when their aspect ratio exceeds a critical value [28, 29, 30], resulting in the formation of plasmoids/flux ropes that drive fast magnetic reconnection to convert magnetic energy into plasma kinetic energy [31, 32]. Eventually, the plasma is expelled from the reconnection layer and the tension in the magnetic field is released as the plasma flows out. The magnetic field lines are then re-extended by the dragging of the black hole to form an anti-parallel configuration, and the current sheet is induced again to repeat the magnetic reconnection process. During each magnetic reconnection, the plasma in the current sheet splits into two parts, one co-rotating with the black hole and the other counter-rotating. The co-rotating part is accelerated while the counter-rotating part is decelerated. Similar to the Penrose process, if the decelerated part has negative energy and is absorbed by the black hole, while the accelerated part escapes, the accelerated part ends up carrying extra energy, thus realizing energy extraction from the black hole. This extra energy comes from the black hole’s rotational energy.
@@
Astronomical observations have confirmed that there are often different scales of magnetic fields around black holes due to the presence of accretion matter or companion stars. For example, the supermassive black hole Sagittarius A* is accompanied by the magnetar SGR J1745-2900 [33, 34, 35, 36], and a strong magnetic field is also present near the event horizon of the black hole in M87* [37]. Therefore, the process of magnetic reconnection of black holes can be expected to occur routinely. In fact, in the BZ mechanism, energy extraction is also realized with the help of the magnetic field around the black hole. The Comisso-Asenjo mechanism is distinguished from the BZ mechanism by its bursty nature and reliance on non-zero particle inertia. Studies by Comisso and Asenjo suggest that this mechanism may be more efficient than the BZ process in certain parameter spaces, making it a promising candidate for energy extraction from black holes.
@@
On the other hand, to test GR and the Kerr hypothesis, it is necessary to construct other non-Kerr rotating black holes. There are usually two ways to construct them, the top-down approach and the bottom-up approach. In the top-down method, one starts from a modified gravitational theory and obtains exact rotating black hole solutions by solving the gravitational field equations. However, solving modified gravitational field equations is generally very challenging due to their complexity. Only in a few cases or under certain approximations (e.g., slow rotation) can exact analytical or numerical rotating black hole solutions be obtained, e.g., the Einstein-dilaton-Gauss-Bonnet black holes [38, 39, 40, 41, 42], the Chern-Simons black holes [43, 44, 45, 46], and the Kerr-Sen black holes [47]. Conversely, the bottom-up method involves constructing parameterized black holes by generalizing the Kerr metric based on considerations like symmetries and regularity, which are independent of any specific gravity theory [48, 49, 50, 51, 52]. These non-Kerr rotating black holes usually contain several deformation parameters introduced to characterize the deviations from the Kerr black hole. Several of them align well with present-day astronomical observations, indicating that their presence cannot be discounted. Consequently, they may offer viable alternatives to the Kerr metric given the current level of accuracy in astronomical observations. Studies of black hole energy extraction via the Comisso-Asenjo mechanism have also been extended to these non-Kerr rotating black holes, such as spinning braneworld black holes [53], the Johannsen-Psaltis metric [54], Kerr-Sen and Kiselev black holes [55], rotating black holes with broken Lorentz symmetry [56], spinning regular black hole [57], spinning hairy black hole [58], Kerr-MOG black holes [59], Kerr-Newman-MOG black holes [60] and Kerr-de Sitter black holes [61]. These studies have demonstrated that deviations from the Kerr black hole significantly impact the energy extraction process.
@@
Inspired by these studies, our main goal in this paper is to study the energy extraction process via the Comisso-Asenjo mechanism for Konoplya-Rezzolla-Zhidenko (KRZ) parameterized black holes [50]. The KRZ metric is proposed to describe generic stationary and axisymmetric black holes. It has several subtle advantages over other parameterized metrics. In the weak-gravity regime, it can be well aligned with Kerr black holes and thus passes the current astronomical observational constraints, while in the near-horizon strong-gravity regime, it needs only a few dominant parameters to sufficiently describe generic deviations from the Kerr black hole. Moreover, the KRZ metric can be mapped to many top-down metrics when the deformation parameters take on some specific values. Due to these merits, it has been utilized to test GR and the Kerr hypothesis using current astronomical data, leading to the examination of constraints on the deformation parameters [62, 63, 64, 65]. This study primarily focuses on the impact of these deformation parameters, which signify deformations near the horizon, on the magnetic reconnection process and subsequent energy extraction. Given that magnetic reconnection occurs in the near-horizon ergoregion, it is anticipated that these deformation parameters will exert notable influences on this process. Future precise astronomical observations, particularly in the strong field region near the horizon, may facilitate the testing of Kerr’s hypothesis by imposing constraints on these deformation parameters through this process.
@@
The paper is organized as follows. In Section II, a brief overview of the KRZ-parameterized black hole is provided, along with an examination of how the deformation parameters impact the ergoregion. Section III delves into a thorough analysis of the circular geodesic motion of the plasma in the equatorial plane. Section IV investigates the extraction of energy from the black hole via the Comisso-Asenjo mechanism. The last section is the summary and conclusions.
@@ II The KRZ parametrized black hole
In Boyer-Lindquist coordinates, the KRZ metric reads [50, 62, 63]
@@
Here $\{\delta_{i}\}(i=1,2,\cdots,6)$ are six deformation parameters introduced to characterize deviations away from the Kerr metric: $\delta_{1}$ corresponds to deformations of $g_{tt}$, $\delta_{2}$ and $\delta_{3}$ correspond to rotational deformations of the metric, $\delta_{4}$ and $\delta_{5}$ correspond to deformations of $g_{rr}$, and $\delta_{6}$ corresponds to deformations of the event horizon. In the limit of all $\delta_{i}\rightarrow 0$, the KRZ metric reproduces the Kerr metric exactly. Note that in the metric the sign of $a$ can be absorbed into $\{\delta_{2},\delta_{3}\}$ and by redefining $t$, so we will only consider $a>0$. Moreover, as we will see later, among the six deformation parameters, $\{\delta_{1},\delta_{2}\}$ are the most important for the physics we are going to study. So, for simplicity we will set $\delta_{3}=\delta_{4}=\delta_{5}=\delta_{6}=0$ in the following and only consider the influences of $\{\delta_{1},\delta_{2}\}$. The constraint on $\{\delta_{1},\delta_{2}\}$ from X-ray observations of the supermassive BH in Ark 564 is [63]
@@
For non-vanishing deformation parameters, the above equation determines the radius of the event horizon as a function of the polar angle $\theta$. The ergoregion, where magnetic reconnection occurs, lies between the event horizon and the ergosphere (the static limit surface) and is defined as
@@
In the equatorial plane $\theta=\pi/2$, the event horizon is at $r=r_{0}$, while the location of the ergosphere $r=r_{\rm sl}$ is only affected by $\{\delta_{1},\delta_{2}\}$. Fig. 1 shows the ergosphere $r=r_{\rm sl}$ in the equatorial plane as a function of $\{\delta_{1},\delta_{2}\}$ for various values of the spin $a$. It can be seen that the deformation parameters $\{\delta_{1},\delta_{2}\}$ have a significant influence on the shape of the ergoregion. From the left panel, it can be seen that a larger $a$ results in a larger ergoregion. And for a fixed $a$, $r_{\rm sl}$ is a monotonically decreasing function of $\delta_{1}$, meaning that negative $\delta_{1}$ enlarges the ergoregion while positive $\delta_{1}$ shrinks it. The effect of $\delta_{2}$ on $r_{\rm sl}$ is more complicated, as shown in the right panel. For fixed and large $a$, $r_{\rm sl}$ is a monotonically increasing function of $\delta_{2}$, meaning that positive $\delta_{2}$ will expand the ergoregion, while negative $\delta_{2}$ will contract it. While for fixed and small $a$, $r_{\rm sl}$ is a convex function of $\delta_{2}$, meaning that large negative $\delta_{2}$ will also expand the ergoregion in this case. Note that if $\delta_{2}$ is sufficiently negative, a higher (but still small) spin $a$ is not necessary to produce a larger ergoregion (e.g., the ergoregion for $a=0.5$ is smaller than that for $a=0.1$ when $\delta_{2}$ is large negative).
@@ III Circular geodesics motion in equatorial plane
where dot means derivative with respect to some affine parameter. From the conservation of the rest mass, ${\cal L}=-1/2$, we have
@@
For circular orbits in the equatorial plane $\theta=\pi/2$, the effective potential should satisfy the following three conditions
@@
The last condition can always be satisfied due to the reflection symmetry of the spacetime with respect to the equatorial plane. With the first two conditions, one can obtain the Keplerian angular velocity $\Omega_{K}\equiv\dot{\phi}/\dot{t}$, the specific energy $E$ and the $z$-component angular momentum $L_{z}$ of the particle
@@
where the sign $+$ stands for co-rotating orbits and -- for counter-rotating orbits. Note that of the six deformation parameters, only $\{\delta_{1},\delta_{2}\}$ enter the three equations above. Given the values of the spin $a$ and the deformation parameters $\{\delta_{1},\delta_{2}\}$, the above equations determine $\Omega_{K},E$ and $L_{z}$ as a function of the radius of the circular orbit. The circular orbit is stable under perturbations if $\partial_{r}^{2}V_{\rm eff}\leq 0$ and $\partial_{\theta}^{2}V_{\rm eff}\leq 0$. A radially stable circular orbit exists from infinity to the innermost (radially) stable circular orbit (ISCO), which is given by the condition
@@
Figure 2: Radius of innermost stable circular orbit as a function of the spin $a$. Here, we fix $\delta_{2}=0$ and study the influence of $\delta_{1}$. The Kerr case $(\delta_{1}=\delta_{2}=0)$ is also shown for comparison. $r^{+}_{\rm ISCO}$ and $r^{-}_{\rm ISCO}$ represent the co-rotating and counter-rotating orbits, respectively. $r_{0}$ is the event horizon and $r_{sl}$ is the ergosphere.
Figure 3: Radius of innermost stable circular orbit as a function of the spin $a$. Here we fix $\delta_{1}=0$ and study the influence of $\delta_{2}$. $r^{+}_{\rm ISCO}$ and $r^{-}_{\rm ISCO}$ represent the co-rotating and counter-rotating orbits, respectively. $r_{0}$ is the event horizon and $r_{sl}$ is the ergosphere.
In Figs. 2 and 3 we plot the radius of the ISCO as a function of the spin $a$ for several typical values of $\{\delta_{1},\delta_{2}\}$. The Kerr case is also shown for comparison. In each figure, we set one of the two parameters $\{\delta_{1},\delta_{2}\}$ to zero and study the influence of the other. From the figures, we can see that as the spin $a$ increases, the radius of the counter-rotating ISCO $r^{-}_{\rm ISCO}$ increases monotonically and is always outside the ergoregion. Therefore, we will focus only on the co-rotating orbits. As the spin $a$ is increased, the radius of the co-rotating ISCO $r^{+}_{\rm ISCO}$ decreases almost monotonically. The only exception is when $\delta_{1}<0$ and $a$ is close to its extreme value $a=1$, where $r^{+}_{\rm ISCO}$ increases with increasing $a$. The co-rotating ISCO will enter the ergoregion when the spin $a$ is high enough to exceed some critical value $a_{c}$. From the figures, one can see that $\delta_{1}$ has little effect on $a_{c}$, while $\delta_{2}$ has a significant effect on $a_{c}$. As $|\delta_{2}|$ is increased, $a_{c}$ increases for $\delta_{2}<0$ and decreases for $\delta_{2}>0$. Since magnetic reconnection is related to orbits within the ergoregion, these results imply that for larger $|\delta_{2}|$, magnetic reconnection can occur for lower black hole spin if $\delta_{2}>0$, while for $\delta_{2}<0$ higher spin is required for magnetic reconnection to occur. Furthermore, $r^{+}_{\rm ISCO}$ does not coincide with the event horizon $r_{0}$ in the extreme limit when $\delta_{1}\neq 0$, and $r^{+}_{\rm ISCO}\neq r^{-}_{\rm ISCO}$ in the non-rotating limit $a=0$ if $\delta_{2}\neq 0$.

### v2

# Energy Extraction via Magnetic Reconnection in Konoplya-Rezzolla-Zhidenko Parametrized Black Holes

@@ Abstract
Recently, Comisso and Asenjo proposed a novel mechanism for harnessing energy from black holes through magnetic reconnection. Our study focuses on exploring the utilization of this mechanism on Konoplya-Rezzolla-Zhidenko (KRZ) parametrized black holes to assess the impact of deformation parameters on energy extraction. Among the various parameters, $\{\delta_{1},\delta_{2}\}$ are identified as the most important factors affecting the physics under consideration. The influence of these two parameters on the ergoregion, circular geodesics in the equatorial plane, and energy extraction from the KRZ black holes through magnetic reconnection is carefully analyzed. Results indicate that deviations from the Kerr metric have a notable influence on the energy extraction process. Particularly, energy extraction is enhanced with more negative ${\delta_{1}}$ or more positive ${\delta_{2}}$ within the range constrained by current astronomical observations, resulting in significantly higher maximum power and efficiency compared to the Kerr model. Moreover, when ${\delta_{2}}$ is sufficiently negative, extracting energy through this mechanism becomes increasingly challenging, necessitating an exceptionally high black hole spin.
@@ I Introduction
The first plausible mechanism for energy extraction from black holes was proposed by Penrose in 1969, now known as the Penrose process [8]. The Penrose process involves the splitting of a particle in the ergoregion into two particles. Measured from infinity, one of these particles may carry negative energy, while the other carries excess positive energy. The extraction of energy from a black hole occurs when the negative energy particle is absorbed by the black hole and the positive energy particle escapes. However, the Penrose process is considered impractical because it requires the two newborn particles to have relative velocities greater than 1/2 the speed of light, an event thought to be rare in real astrophysical processes [9, 10]. Inspired by Penrose’s work, physicists have proposed several alternative mechanisms for harnessing energy from black holes, including superradiant scattering [11], the collisional Penrose process [12], the Blandford-Znajek (BZ) process [13], and the magnetohydrodynamic (MHD) Penrose process [14]. Among these mechanisms, the BZ process is currently regarded as the most promising for elucidating GRBs [15, 16, 17] and relativistic jets in AGNs [18, 19, 20, 21].
@@
Recently, based on an exploratory study [22], a novel mechanism for black hole energy extraction has been proposed by Comisso and Asenjo [23]. In this mechanism, energy extraction is realized by a process of magnetic reconnection in the ergoregion of a rotating black hole. The process begins with the magnetic field around the black hole forming antiparallel configurations near the equatorial plane due to the black hole’s rotation [24, 25, 26, 27]. The jumps in the magnetic field direction induce the formation of current sheets between them, which are destroyed by plasmoid instability when their aspect ratio exceeds a critical value [28, 29, 30], resulting in the formation of plasmoids/flux ropes that drive fast magnetic reconnection to convert magnetic energy into plasma kinetic energy [31, 32]. Eventually, the plasma is expelled from the reconnection layer and the tension in the magnetic field is released as the plasma flows out. The magnetic field lines are then reextended by the dragging of the black hole to form an antiparallel configuration, and the current sheet is induced again to repeat the magnetic reconnection process. During each magnetic reconnection, the plasma in the current sheet splits into two parts, one corotating with the black hole and the other counter-rotating. The corotating part is accelerated while the counter-rotating part is decelerated. Similar to the Penrose process, if the decelerated part has negative energy and is absorbed by the black hole, while the accelerated part escapes, the accelerated part ends up carrying extra energy, thus realizing energy extraction from the black hole. This extra energy comes from the black hole’s rotational energy.
@@
Astronomical observations have confirmed that there are often different scales of magnetic fields around black holes due to the presence of accretion matter or companion stars. For example, the supermassive black hole Sagittarius A* is accompanied by the magnetar SGR J1745-2900 [33, 34, 35, 36], and a strong magnetic field is also present near the event horizon of the black hole in M87* [37]. Therefore, the process of magnetic reconnection of black holes can be expected to occur routinely. In fact, in the BZ mechanism, energy extraction is also realized with the help of the magnetic field around the black hole. The Comisso-Asenjo mechanism is distinguished from the BZ mechanism by its bursty nature and reliance on nonzero particle inertia. Studies by Comisso and Asenjo suggest that this mechanism may be more efficient than the BZ process in certain parameter spaces, making it a promising candidate for energy extraction from black holes.
@@
On the other hand, to test GR and the Kerr hypothesis, it is necessary to construct other non-Kerr rotating black holes. There are usually two ways to construct them, the top-down approach and the bottom-up approach. In the top-down method, one starts from a modified gravitational theory and obtains exact rotating black hole solutions by solving the gravitational field equations. However, solving modified gravitational field equations is generally very challenging due to their complexity. Only in a few cases or under certain approximations (e.g., slow rotation) can exact analytical or numerical rotating black hole solutions be obtained, e.g., the Einstein-dilaton-Gauss-Bonnet black holes [38, 39, 40, 41, 42], the Chern-Simons black holes [43, 44, 45, 46], and the Kerr-Sen black holes [47]. Conversely, the bottom-up method involves constructing parametrized black holes by generalizing the Kerr metric based on considerations like symmetries and regularity, which are independent of any specific gravity theory [48, 49, 50, 51, 52]. These non-Kerr rotating black holes usually contain several deformation parameters introduced to characterize the deviations from the Kerr black hole. Several of them align well with present-day astronomical observations, indicating that their presence cannot be discounted. Consequently, they may offer viable alternatives to the Kerr metric given the current level of accuracy in astronomical observations. Studies of black hole energy extraction via the Comisso-Asenjo mechanism have also been extended to these non-Kerr rotating black holes, such as spinning braneworld black holes [53], the Johannsen-Psaltis metric [54], Kerr-Sen and Kiselev black holes [55], rotating black holes with broken Lorentz symmetry [56], spinning regular black hole [57], spinning hairy black hole [58], Kerr-MOG black holes [59], Kerr-Newman-MOG black holes [60] and Kerr-de Sitter black holes [61]. These studies have demonstrated that deviations from the Kerr black hole significantly impact the energy extraction process.
@@
Inspired by these studies, our main goal in this paper is to study the energy extraction process via the Comisso-Asenjo mechanism for KRZ parametrized black holes [50]. The KRZ metric is proposed to describe generic stationary and axisymmetric black holes. It has several subtle advantages over other parametrized metrics. In the weak-gravity regime, it can be well-aligned with Kerr black holes and thus passes the current astronomical observational constraints, while in the near-horizon strong-gravity regime, it needs only a few dominant parameters to sufficiently describe generic deviations from the Kerr black hole. Moreover, the KRZ metric can be mapped to many top-down metrics when the deformation parameters take on some specific values. Due to these merits, it has been utilized to test GR and the Kerr hypothesis using current astronomical data, leading to the examination of constraints on the deformation parameters [62, 63, 64, 65, 66, 67]. This study primarily focuses on the impact of these deformation parameters, which signify deformations near the horizon, on the magnetic reconnection process and subsequent energy extraction. Given that magnetic reconnection occurs in the near-horizon ergoregion, it is anticipated that these deformation parameters will exert notable influences on this process. Future precise astronomical observations, particularly in the strong field region near the horizon, may facilitate the testing of Kerr’s hypothesis by imposing constraints on these deformation parameters through this process.
@@
The paper is organized as follows. In Sec. II, a brief overview of the KRZ-parametrized black hole is provided, along with an examination of how the deformation parameters impact the ergoregion. Section III delves into a thorough analysis of the circular geodesic motion of the plasma in the equatorial plane. Section IV investigates the extraction of energy from the black hole via the Comisso-Asenjo mechanism. The last section is the summary and conclusions. Throughout this work, we utilize the units $c=G=1$, where $c$ and $G$ are the speed of light in vacuum and the Newton gravitational constant, respectively.
@@ II The KRZ parametrized black hole
In Boyer-Lindquist coordinates, the KRZ metric reads [50, 62, 63],
@@
Note that in the metric, the black hole mass has been set to be $M=1$, so all physical quantities are measured in units of $M$ in the following. Here $\{\delta_{i}\}(i=1,2,\cdots,6)$ are six deformation parameters introduced to characterize deviations away from the Kerr metric: $\delta_{1}$ corresponds to deformations of $g_{tt}$, $\delta_{2}$ and $\delta_{3}$ correspond to rotational deformations of the metric, $\delta_{4}$ and $\delta_{5}$ correspond to deformations of $g_{rr}$, and $\delta_{6}$ corresponds to deformations of the event horizon. In the limit of all $\delta_{i}\rightarrow 0$, the KRZ metric reproduces the Kerr metric exactly. Note that in the metric the sign of $a$ can be absorbed into $\{\delta_{2},\delta_{3}\}$ and by redefining $t$, so we will only consider $a>0$. Moreover, as we see later, among the six deformation parameters, $\{\delta_{1},\delta_{2}\}$ are the most important for the physics we are going to study. So, for simplicity we will set $\delta_{3}=\delta_{4}=\delta_{5}=\delta_{6}=0$ in the following and only consider the influences of $\{\delta_{1},\delta_{2}\}$. The constraint on $\{\delta_{1},\delta_{2}\}$ from x-ray observations of the supermassive BH in Ark 564 is [63]
@@
For nonvanishing deformation parameters, the above equation determines the radius of the event horizon as a function of the polar angle $\theta$. The ergoregion, where magnetic reconnection occurs, lies between the event horizon and the ergosphere (the static limit surface) and is defined as
@@
In the equatorial plane $\theta=\pi/2$, the event horizon is at $r=r_{0}$, while the location of the ergosphere $r=r_{\rm sl}$ is only affected by $\{\delta_{1},\delta_{2}\}$. Figure 1 shows the ergosphere $r=r_{\rm sl}$ in the equatorial plane as a function of $\{\delta_{1},\delta_{2}\}$ for various values of the spin $a$. It can be seen that the deformation parameters $\{\delta_{1},\delta_{2}\}$ have a significant influence on the shape of the ergoregion. From the left panel, it can be seen that a larger $a$ results in a larger ergoregion. And for a fixed $a$, $r_{\rm sl}$ is a monotonically decreasing function of $\delta_{1}$, meaning that negative $\delta_{1}$ enlarges the ergoregion while positive $\delta_{1}$ shrinks it. The effect of $\delta_{2}$ on $r_{\rm sl}$ is more complicated, as shown in the right panel. For fixed and large $a$, $r_{\rm sl}$ is a monotonically increasing function of $\delta_{2}$, meaning that positive $\delta_{2}$ will expand the ergoregion, while negative $\delta_{2}$ will contract it. While for fixed and small $a$, $r_{\rm sl}$ is a convex function of $\delta_{2}$, meaning that large negative $\delta_{2}$ will also expand the ergoregion in this case. Note that if $\delta_{2}$ is sufficiently negative, a higher (but still small) spin $a$ is not necessary to produce a larger ergoregion (e.g., the ergoregion for $a=0.5$ is smaller than that for $a=0.1$ when $\delta_{2}$ is large negative).
@@ III Circular geodesics motion in equatorial plane
where dot means the derivative with respect to some affine parameter. From the conservation of the rest mass, ${\cal L}=-1/2$, we have
@@
For circular orbits in the equatorial plane $\theta=\pi/2$, the effective potential should satisfy the following three conditions:
@@
The last condition can always be satisfied due to the reflection symmetry of the spacetime with respect to the equatorial plane. With the first two conditions, one can obtain the Keplerian angular velocity $\Omega_{K}\equiv\dot{\phi}/\dot{t}$, the specific energy $E$ and the $z$-component angular momentum $L_{z}$ of the particle,
@@
where the sign $+$ stands for corotating orbits and -- for counter-rotating orbits. Note that of the six deformation parameters, only $\{\delta_{1},\delta_{2}\}$ enter the three equations above. Given the values of the spin $a$ and the deformation parameters $\{\delta_{1},\delta_{2}\}$, the above equations determine $\Omega_{K},E$ and $L_{z}$ as a function of the radius of the circular orbit. The circular orbit is stable under perturbations if $\partial_{r}^{2}V_{\rm eff}\leq 0$ and $\partial_{\theta}^{2}V_{\rm eff}\leq 0$. A radially stable circular orbit exists from infinity to the innermost (radially) stable circular orbit (ISCO), which is given by the condition
@@
Figure 2: Radius of innermost stable circular orbit as a function of the spin $a$. Here, we fix $\delta_{2}=0$ and study the influence of $\delta_{1}$. The Kerr case $(\delta_{1}=\delta_{2}=0)$ is also shown for comparison. $r^{+}_{\rm ISCO}$ and $r^{-}_{\rm ISCO}$ represent the corotating and counter-rotating orbits, respectively. $r_{0}$ is the event horizon and $r_{sl}$ is the ergosphere.
Figure 3: Radius of innermost stable circular orbit as a function of the spin $a$. Here we fix $\delta_{1}=0$ and study the influence of $\delta_{2}$. $r^{+}_{\rm ISCO}$ and $r^{-}_{\rm ISCO}$ represent the corotating and counter-rotating orbits, respectively. $r_{0}$ is the event horizon and $r_{sl}$ is the ergosphere.
In Figs. 2 and 3 we plot the radius of the ISCO as a function of the spin $a$ for several typical values of $\{\delta_{1},\delta_{2}\}$. The Kerr case is also shown for comparison. In each figure, we set one of the two parameters $\{\delta_{1},\delta_{2}\}$ to zero and study the influence of the other. From the figures, we can see that as the spin $a$ increases, the radius of the counter-rotating ISCO $r^{-}_{\rm ISCO}$ increases monotonically and is always outside the ergoregion. Therefore, we will focus only on the corotating orbits. As the spin $a$ is increased, the radius of the corotating ISCO $r^{+}_{\rm ISCO}$ decreases almost monotonically. The only exception is when $\delta_{1}<0$ and $a$ is close to its extreme value $a=1$, where $r^{+}_{\rm ISCO}$ increases with increasing $a$. The corotating ISCO will enter the ergoregion when the spin $a$ is high enough to exceed some critical value $a_{c}$. From the figures, one can see that $\delta_{1}$ has little effect on $a_{c}$, while $\delta_{2}$ has a significant effect on $a_{c}$. As $|\delta_{2}|$ is increased, $a_{c}$ increases for $\delta_{2}<0$ and decreases for $\delta_{2}>0$. Since magnetic reconnection is related to orbits within the ergoregion, these results imply that for larger $|\delta_{2}|$, magnetic reconnection can occur for lower black hole spin if $\delta_{2}>0$, while for $\delta_{2}<0$ higher spin is required for magnetic reconnection to occur. Furthermore, $r^{+}_{\rm ISCO}$ does not coincide with the event horizon $r_{0}$ in the extreme limit when $\delta_{1}\neq 0$, and $r^{+}_{\rm ISCO}\neq r^{-}_{\rm ISCO}$ in the nonrotating limit $a=0$ if $\delta_{2}\neq 0$.